OpenStudy (anonymous):

H_E_L_P_! Your test scores in science are 83 and 87. What possible scores can you earn on your next test to have a test average between 85 and 90, inclusive?

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

83 + 87 + x all divided by 3

OpenStudy (danjs):

equals

OpenStudy (danjs):

85

OpenStudy (danjs):

that is the lowest

OpenStudy (danjs):

For the lowest score \[\frac{ 83 + 8 7 + x }{ 3 } = 85\]

OpenStudy (anonymous):

Ok what does inclusive mean ?

OpenStudy (danjs):

multiply both sides by 3 83 + 87 + x = 255 take off 83 and 87 from both sides x = 255 - 83 - 87 = 85 To get an average of 85, you need at least an 85 on the next test

OpenStudy (danjs):

Do the same for the higher average solve (83 + 87 + x)/3 = 90

OpenStudy (danjs):

inclusive means, including 85 and 90 \[85 \le \frac{ 83 + 87 + x }{ 3 }\le 90\] the equals part in the less than or equal to

OpenStudy (anonymous):

oh Gotcha

OpenStudy (danjs):

(83 + 87 + x)/3 = 90 solve that for the upper bound